### v1

# Spectropolarimetric Inversion in Four Dimensions with Deep Learning (SPIn4D):

Abstract
The National Science Foundation’s Daniel K. Inouye Solar Telescope (DKIST) will provide high-resolution, multi-line spectropolarimetric observations that are poised to revolutionize our understanding of the Sun. Given the massive data volume, novel inference techniques are required to unlock its full potential. Here, we provide an overview of our “SPIn4D” project, which aims to develop deep convolutional neural networks (CNNs) for estimating the physical properties of the solar photosphere from DKIST spectropolarimetric observations. We describe the magnetohydrodynamic (MHD) modeling and the Stokes profile synthesis pipeline that produce the simulated output and input data, respectively. These data will be used to train a set of CNNs that can rapidly infer the four-dimensional MHD state vectors by exploiting the spatiotemporally coherent patterns in the Stokes profile time series. Specifically, our radiative MHD model simulates the small-scale dynamo actions that are prevalent in quiet-Sun and plage regions. Six cases with different mean magnetic fields have been conducted; each case covers six solar-hours, totaling TB in data volume. The simulation domain covers at least Mm with km spatial resolution, extending from the upper convection zone up to the temperature minimum region. The outputs are stored at a 40 s cadence. We forward model the Stokes profile of two sets of Fe I lines at and nm, which will be simultaneously observed by DKIST and can better constrain the parameter variations along the line of sight. The MHD model output and the synthetic Stokes profiles are publicly available.
UTF8gbsn
1 Introduction
The solar photosphere, a dynamic layer characterized by diverse plasma and magnetic states, plays an essential role in solar astronomy. Within this layer, processes such as magnetic field emergence, convection, and energy and helicity injection are continuously ongoing and define the structure and evolution of the outer solar atmosphere and the rest of the heliosphere. Understanding the photospheric plasma is therefore key to understanding solar activity. The photosphere is dynamically described by the magnetohydrodynamic (MHD, Priest, 2014) equations, which govern the four-dimensional (4D; three for space plus one for time) evolution of the MHD state vector: the magnetic and velocity vector fields, and the scalar density and pressure (or temperature ) fields. In addition to the MHD equations, radiative transfer plays a dual role in the system. On the one hand, it directly modifies the thermodynamic evolution of the plasma relative to basic MHD via the emission and absorption of photons, while on the other hand, the resulting spectra that reaches the far field (e.g., at ground and space-based telescopes) encodes the state of the system, and thus provides crucial diagnostics of the photospheric plasma (Hubeny & Mihalas, 2015). Notably, the Zeeman and Hanle effects encode the state of magnetic field in polarized radiation, provided it can be properly interpreted (Stenflo, 1994; del Toro Iniesta, 2007).
The Daniel K. Inouye Solar Telescope (DKIST, Rimmele et al., 2020), established by the National Science Foundation, has opened a new era in solar observation with its commencement of scientific operations in 2022. Boasting the world’s largest aperture for solar studies (4 m), DKIST enables unprecedentedly sharp observations (down to 0.03″, or 25 km) with the excellent seeing conditions on Haleakalā and its advanced adaptive optics system. Both the design of the telescope and its instrumentation suite enable measurements of polarized signals with unparalleled accuracy (approximately , Rimmele et al., 2020; de Wijn et al., 2022; Jaeggli et al., 2022; Harrington et al., 2023). Its large aperture means it can match the signal-to-noise ratio (SNR) of deep-exposure quiet Sun observations by the Hinode/Spectro-Polarimeter (Lites et al., 2008) in a fraction of the time ( s versus s). Equipped with cutting-edge instruments like the Visible Spectro-Polarimeter (ViSP, de Wijn et al., 2022), the Diffraction-Limited Near-Infrared Spectro-Polarimeter (DL-NIRSP, Jaeggli et al., 2022), the Cryogenic Near-Infrared Spectropolarimeter (Cryo-NIRSP, Schad et al., 2023), and the Visible Tunable Filter (VTF, Schmidt et al., 2016), DKIST has the capability to explore various solar regions, from the photosphere and chromosphere to the corona. Here we focus on the capabilities of the DL-NIRSP, which currently observes in the near-infrared spectrum using integrating fiber-optic integral-field units (IFU) that enable the simultaneous collection of multi-line Stokes profiles across a continuous field of view (FOV)111DL-NIRSP recently (winter 2023) swapped the fiber-optic IFU for a newly developed image-slicer. The resulting data products will be functionally equivalent for our purposes, but have increased fidelity.. This feature significantly enhances our ability to measure and analyze solar phenomena in detail, particularly the magnetic properties of the small-scale structures.
With the help of the above new facilities, spectropolarimetric observations (typically in the form of wavelength dependent Stokes profiles) for solar physics will significantly expand in the near future. These advanced observations are sensitive to the plasma state variables across the solar atmosphere (del Toro Iniesta & Ruiz Cobo, 2016; Bellot Rubio & Orozco Suárez, 2019, and reference therein), offering a unique tool to probe the solar environment in different heights through sophisticated inversion techniques, whereby the multi-dimensional state of the plasma is inferred from 2D maps of the polarized spectra (Ruiz Cobo & del Toro Iniesta, 1992; Asensio Ramos et al., 2008; Socas-Navarro et al., 2015; Milić & van Noort, 2018; delaCruzRodriguez et al., 2019; Quintero Noda et al., 2021, 2023; Ruiz Cobo et al., 2022; Asensio Ramos & Díaz Baso, 2019; Pastor Yabar et al., 2019; Borrero et al., 2019). For example, initial DKIST/ViSP data measurements of magnetic fields in both quiet and plage regions highlight the utility of spectropolarimetric observation in probing plasma dynamics in otherwise inaccessible environments (Campbell et al., 2023; da Silva Santos et al., 2023; Kuridze et al., 2024). Developing good inversion techniques is thus required to tackle a host of unanswered questions in both solar and plasma physics in general, for instance, probing the mechanisms behind local magnetic dynamos (Vögler & Schüssler, 2007; Stenflo, 2012; Rempel, 2014; Lord et al., 2014), understanding the flux of helicity and energy across the solar surface (Welsch, 2015; Schuck & Antiochos, 2019; Lumme et al., 2019; Thalmann et al., 2021; Liu & Schuck, 2012; Liu et al., 2014, 2023), and unraveling the magnetic foundations necessary for solar eruptions and coronal heating (Antiochos, 1998; Antiochos et al., 1999; Moore et al., 2001; Priest & Forbes, 2002; Sun et al., 2013; Wang et al., 2015; Liu et al., 2016; Chitta et al., 2017; Wyper et al., 2017; Wang et al., 2017; Samanta et al., 2019).
Despite recent advancements in the inversion technology, creating accurate 3D reconstructions of the solar atmosphere still remains a challenge. In particular, knowledge of the physical variables’ distribution on a spatial grid is crucial for the calculation of differential quantities that define key parameters of the system, e.g., the electric current, Lorentz force, helicity and energy flux, and so on. Most current inversion techniques return the physical state of the plasma on an optical depth grid rather than on a spatial grid. Converting to a physical grid along the line of sight (LOS) involves the additional steps of resolving the ambiguity in the inverted magnetic azimuth angle, reconstructing the atmosphere based on dynamic/static assumptions, and defining a vertical offset between each (assumed) independent LOS (Pastor Yabar et al., 2019; Borrero et al., 2019, 2021; Borrero & Pastor Yabar, 2023). The spatiotemporal information of the time-series observations is governed by the MHD equations. Once incorporated, they can help resolve the LOS spatial grid and ambiguity issues while also enhancing the physical accuracy of inversions. However, integrating this information is challenging in traditional 1D inversion techniques based on sophisticated radiative transfer methods. These challenges are exacerbated by the need for powerful computers to process the data and the complex work of creating accurate solar models that match up with scientific laws over time, e.g., Newton’s laws and Maxwell’s equations. Moreover, the arrival of powerful telescopes like DKIST, which can generate about 20 TB of data every day, makes these challenges even greater, testing the limits of our current technology. This huge amount of data highlights the urgent need for innovative, efficient methods to handle and analyze large-volume solar Stokes profile data in the DKIST era.
On the other hand, deep learning (DL), a specialized branch of machine learning (ML), has shown exceptional efficacy in deriving approximate inferences from physics models (Sadowski & Baldi, 2018; Brehmer et al., 2020). The rapid progress of ML in solar physics, especially through DL’s application to Stokes inversion, would significantly enhance our understanding and analytical abilities (Asensio Ramos et al., 2023, and reference therein). Recent studies have highlighted the effectiveness of convolutional neural networks (CNNs, Asensio Ramos & Díaz Baso, 2019; Milić & Gafeira, 2020; Gafeira et al., 2021; Centeno et al., 2022; Chappell & Pereira, 2022; Higgins et al., 2021, 2022; Mistryukova et al., 2023; Rahman et al., 2023, 2024) and other advanced ML algorithms (Bobra & Couvidat, 2015; Bobra & Ilonidis, 2016; Florios et al., 2018; Huang et al., 2018; Nishizuka et al., 2018; Sainz Dalda et al., 2019; Díaz Baso et al., 2022; Vicente Arévalo et al., 2022; Jarolim et al., 2023, 2024b, 2024a; Goodwin et al., 2024) in processing a wide array of solar observations, ranging from the quiet Sun to dynamic solar flares and coronal mass ejections. These models have shown promising results in enhancing accuracy and efficiency, notably outpacing traditional methods in speed without sacrificing analytical complexity. Furthermore, the DL methods can enhance spatial resolution and image denoising, improving the capability of observing small structures (Díaz Baso & Asensio Ramos, 2018; Díaz Baso et al., 2019; Asensio Ramos et al., 2018; Rahman et al., 2020; Song et al., 2022; Eklund, 2023). With the help of time-series observations, it also effectively resolves solar surface flows, with the DeepVel code (Asensio Ramos et al., 2017) demonstrating superior performance compared to conventional methods in analyzing these small-scale structures (Tremblay et al., 2018).
Most of the aforementioned ML models are supervised learning models, requiring a large volume of training data to derive a relationship between the input and target data. Fortunately, modern numerical MHD simulations can now accurately mimic various solar phenomena (Rempel, 2012, 2014; Cheung et al., 2010, 2019; Chen et al., 2017, 2023a, 2023b) and therefore allow the generation of extensive and realistic datasets for DL training. Once trained, DL inversion models are extremely fast to run. For example, in the pioneering work of Asensio Ramos & Díaz Baso (2019), they train 2D CNN models using a radiation-MHD simulation of a sunspot and synthesized Stokes profiles. When applied to observational data from the Hinode/SP, their CNN model inverts a pixel map within ms, orders of magnitude faster than current inversion methods. Additionally, the model could recover the 3D MHD variables on a true spatial grid rather than an optical depth grid, and at roughly half the error compared to the Stokes Inversion based on Response functions (SIR, Ruiz Cobo & del Toro Iniesta, 1992, 2012) code. It is worth mentioning that Asensio Ramos & Díaz Baso (2019) did not treat the azimuthal ambiguity in the magnetic field directly, but instead solved a reduced problem using a coordinate transformation.
To address the challenges outlined above and take advantage of the potential of DL, we launched the “Spectropolarimetric Inversion in Four Dimensions with Deep Learning” (SPIn4D) project222https://ifauh.github.io/SPIN4D/ to train CNN models on radiative MHD (RMHD) simulations with a significantly larger dataset compared to previous efforts. These data are publicly available333http://dtn-itc.ifa.hawaii.edu/spin4d/DR1/. Figure 2 presents a schematic overview of key elements of each phase of the project. Our focus will be on photospheric regions with intermediate field strengths (between quiet sun and plage, or up to a few hundred Gauss when spatially averaged), characterized by relatively simple field geometries. These regions are expected to be prevalent in the initial years of DKIST’s operation. Simulations are carried out using the Max-Planck University-of-Chicago Radiative MHD code (MURaM, Vögler et al., 2005; Rempel et al., 2009; Cheung et al., 2010; Rempel, 2012, 2014) and take the quiet Sun small-scale dynamo simulation (SSD) of Rempel (2014) as the point of departure. Next, we employ forward-modeling of radiative transfer through these simulations using a new version of the SIR code (SIR3D) and the Departure Coefficient aided Stokes Inversion based on Response Functions (DeSIRe, Ruiz Cobo et al., 2022) code to synthesize Stokes profiles for multi-line observations. We have selected the well-studied Fe I lines at nm, nm, nm, and nm due to their significant Landé factors (, , , and , respectively) and their wide-spread use in ground and space-based observations with high Stokes SNR. The third phase involves developing CNN models that aim to accurately correlate time series of observational data, especially those from the DL-NIRSP instrument at DKIST, with precise 4D MHD states. The models will be rigorously trained and evaluated using the data generated in the first two steps. The ability of this model will be compared with the SIR inversion code as a base line. Higher-level variables such the vector velocity and associated Poynting flux across the photosphere may also be estimated directly. The comprehensive workflow and expected results are outlined in Figure 2.
In this paper, we focus on the generation of the training data for the SPIn4D project. The paper is organized as follows: Section 2 delves into the specifics of the MURaM SSD simulation. The process of multi-line synthesis is explored in Section 3. Section 4 addresses simulation artifacts and their treatment. Finally, a summary is provided in Section 5. Additional details on the pipeline are included in the Appendix.
| Modified Field (G)aaThe magnetic field added to the relaxed SSD O16bM atmosphere as the initial condition for our simulations. | Duration (hrs)bbThe physical time of the total simulations for each case. | Size (TB)ccThe total size of the output 3D atmosphere files. | |
| Case 1 | 0 | 5.97 | 12 |
| Case 2 | 6.21 | 12 | |
| Case 3 | ddCase 3 has G. | 6.23 | 12 |
| Case 4 | 6.02 | 12 | |
| Case 5 | 6.04 | 12 | |
| Case 6 | eeCase 6 has G, G, and G added to in three quadrants (see Figure 6). | 6.20 | 49 |
Note. —
2 Solar Atmosphere Simulation
We ran six RMHD simulations covering a variety of photospheric conditions ranging from very quiet Sun to fairly strong plage. Synthetic spectropolarimetric observations are created by running radiative forward models to through the RMHD output (Section 3). The RMHD simulations use the MURaM code (Vögler et al., 2005; Rempel et al., 2009) and take the relaxed solar atmosphere of the small scale dynamo case SSD 016bM from Rempel (2014) as their starting point. Compared to Rempel (2014), the simulation domain was extended by 500 km in the vertical direction above the photosphere and the vertical grid spacing was reduced from to km. This formed the basis of our Case 1, a straightforward continuation of the SSD O16bM simulation. Cases 2, 4, and 5 simulate regions with increasingly stronger average field strength, introduced as an additional uniform magnetic field to each initial condition’s vertical component, , at strengths of G, G, and G, respectively. Case 3 was augmented with an inclined initial magnetic field instead, uniformly set at G across all magnetic components.
The computational domain for Cases 1–5 spanned Mm, with a spatial resolution of km. The horizontal grid size is approximately half of the diffraction limit of DKIST. We saved data output every seconds to match the expected cadence from DL-NIRSP observations. The side boundaries of the domain are periodic, the bottom boundary is open for convective flows as detailed in Rempel (2014), and the top boundary applies a potential magnetic field extrapolation in the ghost layers, along with a semi-transparent boundary condition for hydrodynamic variables, i.e. density, velocity, and internal energy. This condition is designed to be open for outflows and closed for inflows, featuring (anti-)symmetric settings in the ghost layers for inflows and outflows, respectively. Radiative transfer calculations may be carried out in arbitrary directions through the simulated domain, but for the tasks described in this work, we take the LOS direction to be the simulation direction and use these two interchangeably.
In Figure 4, we show representative simulation output toward the end of each run for Cases 1–5: the nm continuum intensity (left), LOS velocity (; center), and LOS magnetic field (; right), each extracted from the layer of optical depth at roughly 5.5 hrs into each run, where is the optical depth at 500 nm. The increasing initial average field strength in each case forms more and stronger field concentrations, as seen in the right column. The greatest addition in Case 5 shows the development of several pores with substantially reduced intensity (left column) associated with the strongest field concentrations (right column), consistent with regions of strong plage on the Sun.
Cases 2–5 each have a net magnetic flux that would mimic coronal hole environments on the actual Sun, with no strong polarities with opposing signs. To account for real-Sun conditions with mixed polarities, we ran an additional Case 6 in a larger domain created by stitching together four of our modified SSD O16bM initial conditions in the two periodic horizontal directions and then adding additional field of strengths G, G, and G to the vertical component in three quadrants; the forth quadrant was left in the initial SSD configuration. To avoid discontinuous field strengths, the vertical field added to each quadrant was first multiplied by a mask function that decreases smoothly to zero at each quadrant boundary,
where and are normalized coordinates in each field region before being stitched into the larger domain.
Representative results for Case 6 are shown in Figure 6 around 5.5 hours into the simulation. Figure 6(d) shows the additional vertical field added to each quadrant in the initial condition. After a few solar hours of relaxation, this simulation showed a mix of strong polarities with interaction of opposite polarity patches at the quadrant boundaries (see Figure 6(c)). Portions of this simulation may therefore represent the boundaries of long–decayed active regions.
Each simulation extends over Mm vertically for all Cases, from a shallow convection zone to the upper photosphere, represented by grid points. The photospheric Fe I lines that we are interested in form in a relatively narrow layer ( Mm) entirely contained within the simulation domain. In Figure 8(a)–(d) we show the response functions of Stokes to the magnetic field magnitudes of these lines calculated using SIR. Based on this, and to optimize storage, we selected a range of layers from the simulation interiors for output (at 40 s cadence), specifically from the grid points to in the vertical direction, spanning an optical depth range from to and demarcated by the blue dashed lines in Figure 8(e). Comparison with the response functions shows that the extracted region adequately encompasses the formation layers of the Fe I lines of interest. Further details are provided in Section 3.
All Cases 1–6 were simulated for about six physical hours and generated a total of TB of output for the 3D MHD variables, including the high-cadence output just described plus occasional snapshots of the full numerical domain. The initial magnetic configuration, total duration, and total output for all cases are summarized in Table 1.
2.1 Initial relaxation, convective turnover time, and statistically independent training data
The addition of vertical fields to the initial conditions of Cases 2–6 represent substantial injections of magnetic energy, and the resulting simulated atmospheres need some time to relax to new steady dynamical states. To illustrate the atmospheric relaxation process, we focus on Case 5, which features the most intense initially added magnetic field at 200 G. Figure 10 presents time series histograms of physical variables computed at surfaces of and in the left and right columns, respectively, i.e., near the maxima of the response functions of the selected Fe I lines (see Figure 8). The histogram density is displayed as a color scale, with the histogram bin values given on the linear-scaled ordinates, and time given the log-scaled abcissas (to highlight the approach toward a quasi-stationary state). Notably, the density distributions exhibit pronounced oscillations attributable to the initial addition of a uniform and the resulting imbalance of total pressure. The histograms for exhibits an initial, rapid reduction of pixels with smaller magnitudes (e.g., less than G), as well as a gradual increase of pixels with larger (e.g., greater than G). Both distributions stabilize after approximately two hours (vertical dashed line).
With the goal of generating a large training dataset suitable for DL, it is important to ensure that each snapshot of training data contains statistically independent information. As the primary spatial feature in our simulations is the solar granulation pattern, we seek to minimize the temporal correlation of this pattern in the final training dataset. To achieve this, we assess the lifetime of granules in 500 nm continuum intensity by calculating the Structural Similarity Index (SSIM) using the skimage.metrics.structural_similarity python package (Wang et al., 2004). SSIM returns a scalar value for the difference between two images, where a value of indicates identical (different) images. For each simulation, we calculate SSIM backwards in time, using the final continuum intensity image as a fixed reference image. The results shown in Figure 12. The curve for each case is offset vertically by to aid legibility. Each SSIM curve approaches after about 10 minutes, indicating significant image variation; such is consistent with a granule’s life time of about minutes. As an initial attempt, we selected a 12-minute cadence to generate the synthesized the Fe I lines for our training dataset. More discussion on this choice can be found in Section 3.
2.2 On the range of simulated solar environments
Cases 1–5 represent a range of conditions found on the Sun, differentiated solely by the strength of the uniform magnetic field added to the base SSD simulation for each Case’s initial condition; Case 6 extends this with a larger FOV and mixed polarity regions. The varied conditions produce corresponding differences in the spatial structure of the resulting granulation pattern. In Figure 14, we quantify those differences by computing, in the surface, the spatial power spectra of the normalized continuum intensity (row 1, a–c), vertical velocity (row 2, d–f), and vertical magnetic field (row 3, g–i). All the curve plots in the log scale in this figure are offset vertically by multiplying a factor of for better visualization. The effect of adding stronger magnetic fields is made more clear by taking the ratio of the power spectra for Cases 2–6 to the SSD case, as shown in the second column (panels b, d, e). Together, the power spectra and their ratios reveal that several spatial scales are either present in all cases or notably arise due to the addition of a magnetic field. These scales are indicated in the Fourier domain by vertical dashed lines and in the physical domain by solid lines drawn on top of a representative subregion for each variable in panels c, f, and i.
Both the continuum intensity and vertical velocity show pronounced peaks at the scale of the granulation pattern, around 1.4 and 1.3 Mm, as seen in panels (a) and (d), respectively. The vertical magnetic field shows a very slight peak for Case 1 only (panel g); the peaks are indicated by the vertical dashed line in the first column. From the power spectral ratios presented in the second column, there is a clear trend toward higher power at higher spatial frequencies relative to the SSD case, and a generally steeper slope for cases with a greater added magnetic field strength. In the continuum intensity, this trend produces the large peak in ratio at a scale of km, which corresponds to the readily apparent small scale intergranular bright points seen in Figure 14(c). There is no peak in ratio in the power spectra of , but instead a saddle point at a scale of km (Figure 14(e)), which is the width of the intergranular lanes.
Turning our attention to the power spectrum of , the only apparent peak occurs for Case 1 and has size Mm, i.e., roughly half of the granule size as measured in either continuum intensity or velocity. Surprisingly, as more flux is added to the simulation the magnetic field becomes more ordered (more distinct local peaks) on larger scales. This somewhat mimics a mesogranular morphology and has two local peaks at and Mm (Figure 14(h,i)). However, this might be influenced by interaction with bottom boundary, since we do not have a deep enough convection layer for the formation of the supergranulation convection pattern (Lord et al., 2014; Lord, 2014).
3 Stokes Profile Synthesis
| Fe I | 6301-6302 Å | 15648-15652 Å |
| Line Range (Å) | 6300.8–6303.3 | 15646.8–15654.8 |
| (mÅ) | 8.945 | 31.376 |
| SIR Data Size (TB) | 5.5 | 5.1 |
| DeSIRe Data Size (TB) | 5.5 | 5.1 |
We calculated synthetic Stokes data for each Case at the statistically independent 12 minute cadence after the initial 2 hours relaxation period, as determined in Section 2. For the Stokes profiles dataset, we selected two Fe I line pairs in the nm and nm range, originating from the deep and upper layers of the photosphere, respectively. These lines, whose response functions are depicted in Figure 8(a)–(d), are pivotal in diagnosing the photospheric magnetic field (see Bellot Rubio & Orozco Suárez, 2019, and reference therein). Employing a multi-line diagnostic approach with these lines will facilitate a comprehensive understanding of the three-dimensional photospheric structure.
| Wave length (Å) | aaEnhancement factor to the van der Waals coefficient. | Exc. Pot bbExcitation potential of the lower level. (eV) | log()ccThe logarithm of the multiplicity of the level times the oscillator strength. | Transition Level | ddThe collisional broadening parameters from the quantum mechanical theory of Barklem et al. (1998). | ddThe collisional broadening parameters from the quantum mechanical theory of Barklem et al. (1998). (cm2) |
| 6301.5012 | 1.0 | 3.654 | -0.718 | 5P 2.0- 5D 2.0 | 0.242 | 2.33543e-14 |
| 6302.4936 | 1.0 | 3.686 | -1.131 | 5P 1.0- 5D 0.0 | 0.239 | 2.38024e-14 |
| 15648.5088 | 1.0 | 5.426 | -0.652 | 7D 1.0- 7D 1.0 | 0.229 | 2.72747e-14 |
| 15652.8809 | 1.0 | 6.246 | -0.050 | 7D 5.0- 7k 4.0eeThis transition level is used for SIR synthesis pipeline, while for the DeSIRe, the level is 7D 5.0- (6D4.5)f2k 4.0. | 0.330 | 4.045e-14 |
Note. — The atomic parameters are written in SIR format (Ruiz Cobo & del Toro Iniesta, 1992).
The output wavelength sampling for the synthesis was chosen to match the capabilities of the DL-NIRSP instrument, as detailed in Jaeggli et al. (2022). For the nm spectral window, the DL-NIRSP has a nominal bandpass (derived from the combination of a narrowband filter and spectral mask) of Å, a spectral sampling rate of approximately mÅ per pixel, and a point spread function (PSF) with a full width at half maximum (FWHM) of about mÅ. Our synthesis was performed at half the DL-NIRSP sample step, about mÅ, covering a range from Å to Å across steps. These settings fully cover both lines and extend well into the continuum on either side.
Similarly, for the nm Fe I line pair, the DL-NIRSP as a nominal bandpass off Å, a spectral sampling rate of mÅ per pixel, and a PSF with FWHM of mÅ. We again synthesized at double the DL-NIRSP spectral sampling, mÅ, covering a wavelength range of Å to Å with steps. The settings for both line pairs are detailed in Table 2.
We performed the spectral synthesis using two different codes. First, for forward synthesis under local thermal equilibrium (LTE) conditions, we used the 3D version of SIR (Ruiz Cobo & del Toro Iniesta, 1992; Asensio Ramos & Díaz Baso, 2019). To align with our dataset and coordinate system, the code underwent several adaptations, with the revised version accessible on GitHub444https://github.com/ifauh/par-sir and details of the modifications described in Appendix A. Acknowledging the significance of non-LTE effects as identified in the research by Smitha et al. (2020, 2021), we also integrated the DeSIRe code (Ruiz Cobo et al., 2022), along with the parallel Python wrapper outlined by Gafeira et al. (2021), into our software pipeline. This approach facilitated the generation of non-LTE Stokes profiles for our chosen lines. The elemental abundance data for SIR source is from Grevesse & Sauval (1998), which is consistent with that used in MURaM simulation. For DeSIRe we use an updated abundance data from Asplund et al. (2009). The atomic parameters for the Fe I lines are detailed in Table 3. The resulting synthesized spectral datasets are illustrated in Figure 16.
Cases 2–5 are characterized by a dominant positive magnetic flux. To account for the equally possible input of negative polarity, we manually flipped the signs of the magnetic field, and calculate a second set of synthetic profiles for each dataset. Note that the plasma evolution under the ideal MHD equations remains unchanged when the sign of all three magnetic field components are inverted. This approach allows for random selection between the original and flipped magnetic fields during training of the machine learning model, effectively minimizing the possible bias from the magnetic field polarity in the resulting neural network model. In total, we generated a collection of TB of Stokes profiles for both SIR and DeSIRe, with and without magnetic field flipping, covering dynamically independent snapshots.
Stokes profiles for five representative locations are shown in Figure 18, corresponding to Case 5 at hrs. The first (b–e) and second (f–i) rows present the Stokes profiles for Fe I nm and nm line pairs from the SIR code. The third (j–m) and fourth (m–q) rows present the difference between the synthesized profiles from SIR and DeSIRe, demonstrating minimal discrepancies (note the scale difference between the upper and lower rows). Figure 20 presents the joint distribution between the SIR and DeSIRe synthesized Stokes profiles for the nm line pair for the same dataset, with the SIR results shown on the abscissa and DeSIRe on the ordinate for each panel. The red dashed curve marks the one-to-one line. From top to bottom, the rows compare wavelengths in the blue wing, line core, and red wing of the nm line, followed by similar plots for the nm line. The Pearson correlation of the Stokes I and V is always greater than , while the linear polarizations (Q and U) show a slightly smaller correlation, . Figure 26 shows the analogous comparison for the nm line pair, in Appendix A. The high correlation suggests that the results from the two synthesis codes are in qualitative agreement. The differences can be owing in part to the LTE versus non-LTE treatment of line formation or the different opacity packages.
Our analysis revealed discrepancies in the continuum opacity as a function of and gas pressure between the SIR and MURaM results. Figure 22 contrasts their values. MURaM’s values are derived from its own lookup tables (Castelli & Kurucz, 2004), while SIR’s are calculated with its PemufromPgT_i module using and . Both use the same element abundance (Grevesse & Sauval, 1998). Notably, in certain optical depth layers where the line is formed (outlined by the colored contours in Figure 22(c)), the opacity from MURaM is slightly larger than that from SIR.
A direction comparison between MURaM and DeSIRe opacity prove to be difficult. The latter uses a different opacity package from the RH code (Uitenbroek, 2001). It is calculated under non-LTE conditions, which does not solely depend on local thermal variables. This is expected to cause some difference in the synthetic profiles shown in Figures 18, 20, and 26. As it is out of the scope of the SPIn4D project, we defer a comprehensive investigation of their differences to future work.
4 Artifacts in MHD Simulation and Solutions
Several features of the MHD simulations may require attention when using the simulation database. In all six cases, significant oscillations occur in the upper regions of the simulated atmosphere, close to the top boundary. These are especially apparent for Case 5, with the strongest magnetic field of G (Figure 24(a)). The root cause for this is that we artificially limited the Alfvén velocity to km s-1 to speed up the computation (Rempel et al., 2009; Rempel, 2017) .This is justified for our primary target, the photosphere, where is low. However, numerical issues may arise in higher atmospheric layers as increases rapidly with decrease density such that the imposed limit becomes inconsistent with the calculation. To address this issue, we adjusted the Courant-Friedrichs-Lewy (CFL) condition and Alfvén velocity limit as necessary to maintain numerical stability. The approach proves to significantly reduce the oscillation, though some residual signals remain. Other solutions include adding proper treatment for a transition region in higher layers. Unfortunately they would lead to more than a tenfold increase in the required computing time, so were not adopted there.
These oscillations are barely detectable under a weak mean magnetic field; specifically, Case 5 is intermittently impacted, whereas Case 4 remains mostly unaffected. As a test, we also simulated a case with a stronger, G added magnetic field, which produced even stronger oscillations (Figure 24(b)). Efforts to stabilize this simulation by reducing the CFL parameter and adjusting velocity parameters were unsuccessful. A potential solution involves expanding the simulation box upwards and incorporating a realistic corona. This would significantly increase computational demands, exceeding the SPIn4D project’s scope. Consequently, we limited our simulations to a maximum of G magnetic field strength.
In both 200 G and 400 G cases, we observed numerous “hot pockets” in the regions close to the top boundaries. In these regions, the code attempts to form a hot layer analogous to the transition region, but the set boundary conditions do not allow it to do so. These hot pockets tend to extend to the lower region as the mean magnetic field increases. Meanwhile, temperatures outside these pockets above Mm rise slightly with a stronger mean magnetic field, though the pockets themselves show little correlation with changes in the magnetic field.
We note that these aforementioned artifacts are not expected to impact our results, as illustrated by Figure 24. The upper panel (a) is for Case 5. The regions of numerical instability produce very high temperatures, exceeding K, saturating the color table in white. These regions produce both the strong oscillations and shocks in the surrounding plasma, but are basically confined to strong magnetic field regions. The magnitude of the oscillations decreases as the density rises in the lower regions of the atmosphere, such that the oscillations are largely confined to layers where . Given the distribution of the response function, as shown in Figure 8(a)–(d), the regions critical for the formation of Fe I lines remain unaffected.
5 Summary
In this work, we provide an overview of the SPIn4D project, which aims at advancing the inversion of spectropolarimetric data through machine learning. We describe the procedures for generating a comprehensive training and test dataset derived from MURaM simulations and the forward synthesis of Stokes profiles. Specifically, we conducted six distinct MURaM SSD simulations, generating a total of 109 TB of photospheric atmosphere data. Additionally, we synthesized Fe I lines at nm and nm for every -minute from the simulations, yielding 21 TB of data in HDF5 format. The simulations required an extensive computational effort, amounting to 10 million CPU hours. We have released the SIR-based synthesized Fe I lines and the corresponding 3D photospheric slabs totaling 13.7 TB, making them accessible to the wider research community for further analysis and study. We have used both SIR and DeSIRe codes to synthesize the Stokes profiles. Their results largely agree; additional study is required to explain their minor differences. We will focus on the SIR results and keep the DeSIRe version available for the community. Our DL model training will use these MURaM simulations. The results and comparisons with the SIR inversion as a baseline will be presented in upcoming work.
This dataset, encompassing photospheric physical variables from both quiet Sun and Plage regions, is poised to bolster the burgeoning field of machine learning within solar physics. Its relevance extends particularly to the early research topics of the DKIST science objectives (Rast et al., 2021). The versatility of the data not only supports the inversion tasks of the SPIn4D project but also potentially facilitates the development of a range of other ML models. The public availability of this dataset ensures that the broader scientific community can leverage it not just for inversion studies but also for other tasks, like advanced super-resolution and disambiguation of the horizontal magnetic field. The large volume can reduce the issue of overfitting when training DL models on a small dataset. Its applicability is not limited to DKIST alone; other solar telescopes can also benefit from the insights derived from this comprehensive dataset and the ML model built on it. In this way, the dataset acts as a critical resource, driving forward the integration of machine learning techniques into solar physics and potentially transforming observational strategies and data analysis methodologies across multiple platforms.
Appendix A Changes in SIR3D and Data Organizing
To enhance the handling of half-integer quantum levels for the Fe I nm lines, new atomic state symbols have been updated in the src/interface.f90 module of the 3D version of SIR. The orbital angular momentum range has been expanded from six integer levels to include 13 integer and 11 half-integer levels. The element abundance data was updated in both src/interface.f90 and src/leyendo.f. We have enhanced the selection of optical depth ranges for forward synthesis by introducing new parameters in synth/model.py: clip_tau, clip_tau_min, and clip_tau_max. To address precision issues in the upper layers of MURaM simulations that cause duplication at very small optical depths, line synthesis is now constrained to layers within . We modified the SIR3D code to support a new ’MURAM’ atmosphere that uses solar Cartesian coordinates to match the DL-NIRSP axis ordering and the conventions used in the Python package for the solar community (SunPy). MURaM’s SSD simulation coordinates (, , ) correspond to the SIR3D’s coordinates (, , ) and to solar Cartesian coordinates (, , ), thus a transpose(1,0,2) transformation is applied to all memmap objects in synth/multiprocessing.py to align with the solar Cartesian coordinate system. See our project website for a detailed explanation of these changes.
On the SPIn4D project website, aur datasets are stored within corresponding case directories, whose names are characterized by the mean magnetic field strength. The directories are labeled as follows: SPIN4D_SSD, SPIN4D_SSD_50G, SPIN4D_SSD_50G_V, SPIN4D_SSD_100G, SPIN4D_SSD_200G, and SPIN4D_SSD_Large, corresponding to Cases 1 through 6. For instance, the file subdomain_0.051405 in the directory SPIN4D_SSD_50G corresponds to the mass density data from the simulation output at time index "051405", where "0" is the variable index. Variable indices from 0 to 11 represent mass density (), velocities (, , ), internal energy (), magnetic fields (, , ), temperature (), pressure (), electron number density (), and optical depth (), the shift of the vector components reflecting the coordinate transformations between SIR and MURaM aforementioned. Furthermore, stokes-051405-6302.h5 corresponds to the Fe I 630 nm line at the corresponding output time index. Detailed guidance on accessing and interpreting the data is provided in our online tutorial555https://github.com/ifauh/spin4d-data.
Appendix B Comparison of Synthesized Lines
Statistical comparisons of the synthesized Stokes profiles for the nm line from SIR and DeSIRe are detailed in Figure 26. These discrepancies of the Stokes and are more obvious than those observed in the nm lines shown in Figure 20. Specifically, for the nm line, DeSIRe exhibits greater slopes for Stokes and compared to the identity relation. Despite these differences, the Pearson correlation coefficient remains above for all line positions and all Stokes components, though it shows a weaker correlation than the nm lines, whose smallest correlation is .

### v3

@@
-Spectropolarimetric Inversion in Four Dimensions with Deep Learning (SPIn4D):
+Spectropolarimetric Inversion in Four Dimensions with Deep Learning (SPIn4D): I. Overview, Magnetohydrodynamic Modeling, and Stokes Profile Synthesis
 Abstract
@@
 We forward model the Stokes profile of two sets of Fe I lines at and nm, which will be simultaneously observed by DKIST and can better constrain the parameter variations along the line of sight.
-The MHD model output and the synthetic Stokes profiles are publicly available.
-UTF8gbsn
+The MHD model output and the synthetic Stokes profiles are publicly available, with 13.7 TB in the initial release.
 1 Introduction
@@
 For example, initial DKIST/ViSP data measurements of magnetic fields in both quiet and plage regions highlight the utility of spectropolarimetric observation in probing plasma dynamics in otherwise inaccessible environments (Campbell et al., 2023; da Silva Santos et al., 2023; Kuridze et al., 2024).
-Developing good inversion techniques is thus required to tackle a host of unanswered questions in both solar and plasma physics in general, for instance, probing the mechanisms behind local magnetic dynamos (Vögler & Schüssler, 2007; Stenflo, 2012; Rempel, 2014; Lord et al., 2014), understanding the flux of helicity and energy across the solar surface (Welsch, 2015; Schuck & Antiochos, 2019; Lumme et al., 2019; Thalmann et al., 2021; Liu & Schuck, 2012; Liu et al., 2014, 2023), and unraveling the magnetic foundations necessary for solar eruptions and coronal heating (Antiochos, 1998; Antiochos et al., 1999; Moore et al., 2001; Priest & Forbes, 2002; Sun et al., 2013; Wang et al., 2015; Liu et al., 2016; Chitta et al., 2017; Wyper et al., 2017; Wang et al., 2017; Samanta et al., 2019).
+Developing good inversion techniques is thus required to tackle a host of unanswered questions in both solar and plasma physics in general, for instance, probing the mechanisms behind local magnetic dynamos (Vögler & Schüssler, 2007; Stenflo, 2012; Rempel, 2014; Lord et al., 2014), understanding the flux of helicity and energy across the solar surface (Welsch, 2015; Schuck & Antiochos, 2019; Lumme et al., 2019; Thalmann et al., 2021; Liu & Schuck, 2012a; Liu et al., 2014, 2023), and unraveling the magnetic foundations necessary for solar eruptions and coronal heating (Antiochos, 1998; Antiochos et al., 1999; Moore et al., 2001; Priest & Forbes, 2002; Sun et al., 2013; Wang et al., 2015; Liu et al., 2016; Chitta et al., 2017; Wyper et al., 2017; Wang et al., 2017; Samanta et al., 2019).
 Despite recent advancements in the inversion technology, creating accurate 3D reconstructions of the solar atmosphere still remains a challenge.
@@
 It is worth mentioning that Asensio Ramos & Díaz Baso (2019) did not treat the azimuthal ambiguity in the magnetic field directly, but instead solved a reduced problem using a coordinate transformation.
-To address the challenges outlined above and take advantage of the potential of DL, we launched the “Spectropolarimetric Inversion in Four Dimensions with Deep Learning” (SPIn4D) project222https://ifauh.github.io/SPIN4D/ to train CNN models on radiative MHD (RMHD) simulations with a significantly larger dataset compared to previous efforts.
-These data are publicly available333http://dtn-itc.ifa.hawaii.edu/spin4d/DR1/.
+To address the challenges outlined above and take advantage of the potential of DL, we launched the “Spectropolarimetric Inversion in Four Dimensions with Deep Learning” (SPIn4D) project222https://ifauh.github.io/SPIN4D/.
+to train CNN models on radiative MHD (RMHD) simulations with a significantly larger dataset, i.e., larger domain size, longer evolution, and more cases with a variety of mean magnetic field strengths, compared to previous efforts.
+As discussed below, this provides more statistically independent snapshots useful for training.
+Detailed information about the data, including access methods, are publicly available333Several data access methods are offered for Data Release 1 at http://dtn-itc.ifa.hawaii.edu/spin4d/DR1/..
 Figure 2 presents a schematic overview of key elements of each phase of the project.
@@
 Simulations are carried out using the Max-Planck University-of-Chicago Radiative MHD code (MURaM, Vögler et al., 2005; Rempel et al., 2009; Cheung et al., 2010; Rempel, 2012, 2014) and take the quiet Sun small-scale dynamo simulation (SSD) of Rempel (2014) as the point of departure.
+The fact that these MURaM simulations use gray radiative transfer is not expected to impact the results here.
+As both the training and the evaluation steps use synthetic data, the ML models should simply learn the mapping between the self-consistent inputs and outputs, without regards to the detailed physics.
+When applying to real observations, however, we do expect ML models trained on more realistic, non-gray radiative transfer simulations (Vögler, 2004) to perform better, at the expense of increased computational costs.
 Next, we employ forward-modeling of radiative transfer through these simulations using a new version of the SIR code (SIR3D) and the Departure Coefficient aided Stokes Inversion based on Response Functions (DeSIRe, Ruiz Cobo et al., 2022) code to synthesize Stokes profiles for multi-line observations.
@@
 The ability of this model will be compared with the SIR inversion code as a base line.
-Higher-level variables such the vector velocity and associated Poynting flux across the photosphere may also be estimated directly.
-The comprehensive workflow and expected results are outlined in Figure 2.
+Higher-level variables such the vector velocity and associated Poynting flux across the photosphere may be estimated either based on the inversion results as in conventional methods (e.g., Liu & Schuck, 2012b; Kazachenko et al., 2014), or ultimately directly from our ML models that encode the temporal information with a time-series of Stokes profiles as input.
 In this paper, we focus on the generation of the training data for the SPIn4D project.
@@
 Further details are provided in Section 3.
-All Cases 1–6 were simulated for about six physical hours and generated a total of TB of output for the 3D MHD variables, including the high-cadence output just described plus occasional snapshots of the full numerical domain.
+All Cases 1–6 were simulated for about six physical hours (including the initial two hours of relaxation) and generated a total of TB of output for the 3D MHD variables, including the high-cadence output just described plus occasional snapshots of the full numerical domain.
 The initial magnetic configuration, total duration, and total output for all cases are summarized in Table 1.
@@
 Unfortunately they would lead to more than a tenfold increase in the required computing time, so were not adopted there.
-These oscillations are barely detectable under a weak mean magnetic field; specifically, Case 5 is intermittently impacted, whereas Case 4 remains mostly unaffected.
+It is worth noting that such oscillations do not appear in Rempel (2014), and they are barely detectable for the cases with weaker mean magnetic field.
+Specifically, Case 5 is intermittently impacted, whereas Case 4 remains mostly unaffected.
 As a test, we also simulated a case with a stronger, G added magnetic field, which produced even stronger oscillations (Figure 24(b)).
+The problem arises from both stronger magnetic fields and the more extended vertical domain.
+The combination leads to large values exceeding the imposed limit.
 Efforts to stabilize this simulation by reducing the CFL parameter and adjusting velocity parameters were unsuccessful.
@@
 These regions produce both the strong oscillations and shocks in the surrounding plasma, but are basically confined to strong magnetic field regions.
+Since the MURaM code uses a numerical diffusivity scheme that can significantly enhance diffusivities in regions where monotonicity changes (see Sect.
+2 in Rempel, 2014) where needed and keep them low everywhere else.
+This sufficiently decays and prevents the spread of these oscillations to the lower and higher atmosphere, mainly trapped around Mm where they are generated.
 The magnitude of the oscillations decreases as the density rises in the lower regions of the atmosphere, such that the oscillations are largely confined to layers where .
@@
 To address precision issues in the upper layers of MURaM simulations that cause duplication at very small optical depths, line synthesis is now constrained to layers within .
-We modified the SIR3D code to support a new ’MURAM’ atmosphere that uses solar Cartesian coordinates to match the DL-NIRSP axis ordering and the conventions used in the Python package for the solar community (SunPy).
+We modified the SIR3D code to support a new ‘MURAM’ atmosphere that uses solar Cartesian coordinates to match the DL-NIRSP axis ordering and the conventions used in the Python package for the solar community (SunPy).
 MURaM’s SSD simulation coordinates (, , ) correspond to the SIR3D’s coordinates (, , ) and to solar Cartesian coordinates (, , ), thus a transpose(1,0,2) transformation is applied to all memmap objects in synth/multiprocessing.py to align with the solar Cartesian coordinate system.
 See our project website for a detailed explanation of these changes.
-On the SPIn4D project website, aur datasets are stored within corresponding case directories, whose names are characterized by the mean magnetic field strength.
+On the SPIn4D project website, our datasets are stored within corresponding case directories, whose names are characterized by the mean magnetic field strength.
 The directories are labeled as follows: SPIN4D_SSD, SPIN4D_SSD_50G, SPIN4D_SSD_50G_V, SPIN4D_SSD_100G, SPIN4D_SSD_200G, and SPIN4D_SSD_Large, corresponding to Cases 1 through 6.
@@
 Detailed guidance on accessing and interpreting the data is provided in our online tutorial555https://github.com/ifauh/spin4d-data.
+It is also available on Zenodo https://doi.org/10.5281/zenodo.13879854 (catalog doi:10.5281/zenodo.13879854).
 Appendix B Comparison of Synthesized Lines